Parth (parthkohli):

Math is so scary.\[\text{irrational}^{\text{unreal} \times {\text{irrational}}} = \text{integer} \]That is really scary(Euler's Identity). How can we even have something like this?

Parth (parthkohli):

How can we raise an imaginary power?

Parth (parthkohli):

To give context, \(e^{i\pi} = -1\)

OpenStudy (experimentx):

use Euler's identity ...

OpenStudy (anonymous):

Do you search the derivation of Euler's Identity?

Parth (parthkohli):

Of course, that is Euler's Identity. Exactly, @Spacelimbus Just the intuition.

OpenStudy (anonymous):

Well, I doubt that there is any intuition to be honest, it's a lot of playing around with series and expansion.

Parth (parthkohli):

Anything raised to \(i\) is pretty unreasonable, isn't it?

OpenStudy (anonymous):

If it would be intuitive, we wouldn't be so fascinated about it right?

Parth (parthkohli):

Exactly.

OpenStudy (experimentx):

well ... i admit this i find myself unintuitive. http://www.wolframalpha.com/input/?i=i^i

Parth (parthkohli):

It's surprising how we get no transcendental number. We could still have expected an irrational number, but integer is ridiculous!

OpenStudy (anonymous):

what does "unreal" mean?

Parth (parthkohli):

Imaginary

OpenStudy (anonymous):

ooooooh ok

OpenStudy (anonymous):

like \(e^{\pi i}\) for example

Parth (parthkohli):

@satellite73 Lol that is my second reply.

OpenStudy (anonymous):

Must things about \(i\) seem very unnatural to me at first, I was fascinated by that already: \[ \frac{1}{i} =-i \]

Parth (parthkohli):

That is exactly what I asked. How does this even happen?

OpenStudy (anonymous):

what you have to wrap your head around is what it means to use exponential notation

OpenStudy (anonymous):

you know that \(3^2\) means to multiply 3 by itself, but that is not the meaning of \(3^i\) or even \(3^{\sqrt{2}}\)

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

nice link @experimentX

OpenStudy (anonymous):

once you define \[b^x\] as \[b^x=e^{x\log(b)}\] then you can make more sense out of exponentials, assuming of course that you are familiar with the log

Parth (parthkohli):

Of course I do, continue.

OpenStudy (anonymous):

i was done actually

Parth (parthkohli):

Oh. Let me see.

OpenStudy (anonymous):

but as an example, how would you make sense out of a number like \(2^i\)?

Parth (parthkohli):

\[2 =b, i = x \Longrightarrow e^{i \ln(2)} \]

OpenStudy (anonymous):

yes

Parth (parthkohli):

But I still find it very surprising to get an integer :P

OpenStudy (anonymous):

which is a perfectly good complex number if you want to write it in standard form \(a+bi\) you can write is as \(\cos(\ln(2))+i\sin(\ln(2))\)

OpenStudy (anonymous):

now how about \(e^{\pi i}\) ?

Parth (parthkohli):

Right.

Parth (parthkohli):

Wait a sec, thanks.

OpenStudy (anonymous):

actually i have to run, but i'll be back later.

Parth (parthkohli):

\[ a^b = e^{b \ln(a)} \Longrightarrow e^{i \pi} = e^{i\pi \ln(e)} \]

Parth (parthkohli):

\[\ln(e) = 1 \]

Parth (parthkohli):

Just simplifies to\[ e^{i \pi}\]:/

OpenStudy (anonymous):

\[ \ln e=1 \] But I guess what @satellite73 suggest was using the formula next. \[ \large e^{i \phi }= \text{cis}(\phi)=\cos \phi + i \sin \phi \]

Parth (parthkohli):

That's a whole new thing that I learnt today. What is \(\text{cis}\)? Wait. Must post that as a new question.

Parth (parthkohli):

I don't need the derivation, after all.

OpenStudy (anonymous):

cis is just a lazy notation for cos(x) +isin(x)

OpenStudy (anonymous):

reads like cosines imaginary sinus.

Parth (parthkohli):

I'd do that for sure! (By the way, I am thinking about another problem).

OpenStudy (anonymous):

often \[ \Large re^{i \varphi} \] is just defined as : \[ \Large re^{i \varphi}=r(\cos \varphi +i \sin \varphi)= r\text{cis}\varphi \]

OpenStudy (anonymous):

oh ok.